### 26 Feb 2024 v1

@@ 4 Tight Bounds for Patterns on $n\times n$ Squares
In this section, we prove tight bounds on the tile complexity of self-assembling 2-colored patterns on the surfaces of $n\times n$ squares for almost all positive integers $n$.
@@
For almost all positive integers $n$, the tile complexity of weakly self-assembling 2-colored patterns on a $n\times n$ squares in the aTAM is $\Theta\left(\frac{n^{2}}{\log n}\right)$.
@@
For almost all positive integers $n$, the tile complexity of self-assembling a $2$-color pattern on an $N\times N$ square in the aTAM is $\Omega\left(\frac{n^{2}}{\log n}\right)$.
The proof of Lemma 1 is a straight-forward information-theoretic argument but we include it for the sake of completeness.
@@
Let $\mathcal{T}=(T,\sigma,\tau)$ be a singly-seeded TAS such that $\tau\in\mathbb{Z}^{+}$ is a fixed constant, the strength of every glue of $T$ is bounded by $\tau$ and $B\subseteq T$ and assume $\mathcal{T}$ weakly self-assembles a 2-color pattern on an $n\times n$ square in the aTAM. In other words, every domain of $\mathcal{T}$ is an $n\times n$ square, and we may assume without loss of generality that the type of tile placed at every point therein is either “white” or “black”. Let $B\subseteq T$ be the subset of tiles of $T$ whose colors are “black”.
Let $n\in\mathbb{Z}^{+}$ and $w=w_{n^{2}-1}\cdots w_{0}\in\{0,1\}^{n^{2}}$. Define the pattern corresponding to $w$ as the set $P_{w}\subseteq\{0,\ldots,n-1\}^{2}$ such that for all $1\leq i<n^{2}$, $\left(i\mod n,\left\lfloor\frac{i}{n}\right\rfloor\right)\in P_{w}$ if and only if $w_{i}=1$. Let $P_{w}\subseteq\{0,\ldots,n-1\}^{2}$ and assume $\mathcal{T}_{w}$ weakly self-assembles $P_{w}$.It is easy to see that for every $n\in\mathbb{Z}^{+}$ and $w\in\{0,1\}^{n^{2}}$, there exists a TAS $\mathcal{T}_{w}=\left(T,\sigma,\tau\right)$ such that $\left|T\right|=O\left(n^{2}\right)$.
Going forward, let $n\in\mathbb{Z}^{+}$ and $w\in\{0,1\}^{n^{2}}$ be arbitrary and suppose $\mathcal{T}_{w}$ is the corresponding TAS that weakly self-assembles the pattern corresponding to $w$.
Note that $\mathcal{T}_{w}$ has $4\left|T_{w}\right|$ glues, each strength is bounded by $\tau$, which is a fixed constant, and every tile is either in $B$ or not. This means $\mathcal{T}$ can be represented using $O\left(\left|T_{w}\right|\log\left|T_{w}\right|\right)$ total bits. Let $\left\langle\mathcal{T}_{w}\right\rangle$ be such a representation of $\mathcal{T}_{w}$.
Let $w\in\{0,1\}^{*}$, and $U$ be a fixed universal Turing machine. The Kolomogorov complexity of $w$ is: $K_{U}(w)=\min\left\{\left|\pi\right|\mid U(\pi)=w\right\}$. In other words, $K_{U}(w)$ is the size of the smallest program that when simulated on $U$ outputs $w$. Let $m$ be a non-negative integer and $\varepsilon>0$ be a fixed real constant. The number of binary strings of length less than $m-\varepsilon m$ is at most $1+2+4+\cdots+2^{\lfloor m-\varepsilon m\rfloor-1}=2^{\lfloor m-\varepsilon m\rfloor}-1<2^{\lfloor m-\varepsilon n\rfloor}<2^{m-\varepsilon m+1}$. Define $A_{m,\varepsilon}=\left\{w\in\{0,1\}^{m}\mid K_{U}(w)\geq(1-\varepsilon)m\right\}$. Note that
$\frac{\left|A_{m,\varepsilon}\right|}{2^{m}}\geq\frac{2^{m}-2^{m-\varepsilon m+1}}{2^{m}}=1-\frac{2^{m-\varepsilon m+1}}{2^{m}}=1-\frac{1}{2^{\varepsilon m-1}}.$
Thus, we have
$\lim_{m\rightarrow\infty}{\frac{\left|A_{m,\varepsilon}\right|}{2^{m}}}=1,$
which means that if $\varepsilon>0$ is fixed, then for almost all strings, $w\in\{0,1\}^{m}$, $(1-\varepsilon)n<K_{U}(w)$.
There exists a fixed program $\pi_{SA}$ that takes as input $\langle\mathcal{T}_{w}\rangle$, simulates it, and outputs the string $w$ that corresponds to the pattern $P_{w}$ that $\mathcal{T}_{w}$ self-assembles. Then, for almost all strings $w\in\{0,1\}^{n^{2}}$,
$(1-\varepsilon)|w|<K_{U}(w)<C_{1}\left(\left|\pi_{SA}\right|+\left|T_{w}\right|\log\left|T_{w}\right|\right)<C_{2}\left|T_{w}\right|\log|w|.$
It follows that $\left|T_{w}\right|=\Omega\left(\frac{|w|}{\log|w|}\right)=\Omega\left(\frac{n^{2}}{\log n^{2}}\right)=\Omega\left(\frac{n^{2}}{\log n}\right)$.
Since $n$ and $w$ were arbitrary, the lemma follows.
To prove the lower bound for Theorem 4.1, we prove the following, which is a stronger result that applies to all positive integers $N$.
Lemma 2
For all positive integers $N$, for every pattern $P\subseteq\{0,\ldots,N-1\}^{2}$, there exists an aTAM system $\mathcal{T}=(T,\sigma,1)$ such that $|\sigma|=1$, $|T|=O\left(\frac{N^{2}}{\log N}\right)$ and $\mathcal{T}$ weakly self-assembles $P$.
Proof
@@
We proceed by construction. Let $N\in\mathbb{Z}^{+}$ be the dimensions of the square and $P$ be the $N\times N$ pattern of black and white pixels to weakly self-assemble on the square. Our construction will yield a system $\mathcal{T}=(T,\sigma,1)$ that self-assembles an $N\times N$ square on which $P$ is formed by the black and white tiles of $T$. The tile set $T$ will be composed of two subsets, $T_{s}$ whose tiles form the skeleton, and $T_{r}$ whose tiles form the ribs. We first explain the formation of the skeleton, then that of the ribs. Figure 4 shows a high-level depiction. Throughout our discussion, we will use the notation $P(x,y)$ to denote the color (black or white) of the pattern $P$ at location $(x,y)$.
(b) The square once the ribs of the skeleton have filled in (blue growing left, yellow growing right)..
Figure 4: A schematic example of the construction of the proof of Theorem 2. Instead of showing the black and white colors corresponding to the pattern, we color the tiles to show the pieces of the construction to which they belong.
Skeleton
@@
The seed is part of the skeleton and is placed at location $(\lfloor\log N\rfloor,0)$ and is given the same color as $P(\lfloor\log N\rfloor,0)$. Since a vertical column of the skeleton has width one, and the ribs growing off of each side have length $\lfloor\log N\rfloor$, the width of a pair of ribs and its skeleton column (which we will call a rib-pair) is $12\lfloor\log N\rfloor+$ 1. Dividing the full width $N$ by the width of a rib-pair, and taking the floor, gives the number of full rib-pairs that will fit. Let $f=\lfloor\frac{N}{2\log N+1}\rfloor$ be this number. Let $r=N\mod(2\lfloor\log N\rfloor+1)$ be the remaining width after the last full rib-pair. If $r<\lfloor\log N\rfloor+1$, then a column of the skeleton grows up immediately to the right of the last full rib-pair, and its ribs are of length $r-1$ and grow to the right. If $r\geq\lfloor\log N\rfloor+1$, then the last skeleton column grows upward $\lfloor\log N\rfloor$ positions to the right of the last full rib-pair and has full-length ribs (i.e., $\lfloor\log N\rfloor$) that grow to its left and ribs of length $r-(\lfloor\log N\rfloor+1)$ grow to its right. In the first case, the row of the skeleton that forms the bottom row of the square extends from the seed to $x$-coordinate $f(2\lfloor\log N\rfloor+1)+1$. In the second case, that row extends from the seed to $x$-coordinate $f(2\lfloor\log N\rfloor+1)+\lfloor\log N\rfloor+1$. The tiles of that row are hard-coded and there are $O(N)$ of them. Starting with the seed and then occurring at every $12\lfloor\log N\rfloor+$ 1 locations of the bottom row, a hard-coded set of tiles grows a column of height $N-1$. This row and set of columns are the full skeleton. The number of tile types is $O(N)$ for the row and $O(N)$ for each of the $O\left(\frac{N}{\log N}\right)$ columns, for a total of $O(N)+O\left(\frac{N^{2}}{\log N}\right)=O\left(\frac{N^{2}}{\log N}\right)$ tile types. Note that each skeleton tile type is given the color of the corresponding location in the pattern $P$.
Ribs
@@
From the east and west sides of each location on the columns of the skeleton, ribs grow. Each rib is composed of $\lfloor\log N\rfloor$ tiles (except the ribs growing from the easternmost column, which may be shorter). Since there are two possible colors for each of the $\lfloor\log N\rfloor$ locations of a rib, there are a maximum of $2^{\lfloor\log N\rfloor}\leq N$ possible color patterns for any rib to match the corresponding locations in $P$. (Note that we will discuss the construction of the tiles for ribs that grow to the east, and for ribs that grow to the west the directions are simply reversed.) For any given rib $r$, let the portion of $P$ corresponding to the locations of $r$ be represented by the binary string of length $\lfloor\log N\rfloor$ where each black location is represented by a 00, and each white by a $1$. For example, for a rib $r$ of length $5$ growing eastward from a column, if the corresponding locations of $P$ are “black, black, white, black, white”, then the binary string will be “$00101$”. For each possible binary string $b$ of length $\lfloor\log N\rfloor$, i.e. $b\in\{0,1\}^{\lfloor\log N\rfloor}$, a unique tile type, $t_{b}$, is made. with the glue $b$ on its west side and the glue $b[1:]$ (i.e. $b$ with its left bit truncated) on its east side. This tile type is given the color corresponding to the first bit of $b$. Additionally, for each skeleton column tile from which a rib should grow to the east with pattern $b$, the glue $b$ will be on its east side, allowing $t_{b}$ to attach. This results in the creation of a maximum of $N$ unique tile types (and there will be another $N$ for the first tiles of each westward growing rib). Recall that the tile types for the skeleton were already accounted for and each is hard-coded so that the placement of these glues does not require any new tile types for the skeleton.
@@
Now, the process is repeated for each binary string from length $b-1$ to $1$, with the color of each tile being set to the value of the first remaining bit. Each iteration requires half as many tile types to be created as the previous, i.e. $2^{\lfloor\log N\rfloor-1}$, then $2^{\lfloor\log N\rfloor-2}$, $\ldots$, $2$. Intuitively, each rib position has glues that encode their bit value in the pattern and the portion of the pattern that must be extended outward from them, away from the skeleton. Therefore, for the last position on the tip of each rib, there are exactly 2 choices, white or black, and so all ribs share from a set of two tile types made specially for the ends of ribs. For the tile types of ribs that grow to the east, the total summation is $\Sigma^{\log N-1}_{x=0}{2^{\log N-x}}=2N-2=O(N)$. Accounting for the additional tile types needed for west growing ribs, the full tile complexity of the ribs is $O(N)$.
@@
Correctness of construction
The system $\mathcal{T}$ designed to weakly self-assemble $P$, as discussed, has a seed of a single tile, and since all tile attachments require forming a bond with a single neighbor, the temperature of the system can be $\tau=1$. Our prior analysis shows that the tile complexity is correct at $O\left(\frac{N^{2}}{\log N}\right)$, and showing that $\mathcal{T}$ weakly self-assembles $P$ is trivial since (1) the tiles of the skeleton are specifically hard-coded to be colored for their corresponding locations in $P$, and (2) for each possible pattern corresponding to a rib there is a hard-coded set of rib tiles that match that pattern and grow from the skeleton into that location. Thus, $P$ is formed and Lemma 2 is proved, and with both Lemmas 1 and 2, Theorem 4.1 is proved.
@@ 5 Repeated Patterns
In this section we discuss repeated patterns and show constructions of how to repeat any pattern.
@@
We define a pattern class that given a 2-colored pattern $p_{i}$ and a $m$ will create new pattern $p_{0}$ that tiles $p_{i}$ $m$ times vertically and horizontally forming a grid of the pattern.
$GRID\_REPEAT:(p_{i},m)\rightarrow p_{o}$ where
  • $m\in\mathbb{N}$
  • $p_{i}\in SQPATS_{n}$
  • $p_{o}\in SQPATS_{n*m}$
  • $p_{o}$ is constructed by making an $m\times m$ grid of $p_{i}$ patterns
@@
For all $n,m\in\mathbb{N}$, and $P\in SQPATS_{n}$ there exists an aTAM system $\mathcal{T}=(T,\sigma,2)$ such that $|\sigma|=1$, $|T|=O(\frac{n^{2}}{\log{n}}+\log{mn})$ and $\mathcal{T}$ strictly self-assembles $GRID\_REPEAT(P,m)$.
Proof
We will prove the theorem by presenting a construction for $GRID\_REPEAT$ and showing that all produced aTAMs have $O(\frac{n^{2}}{\log{n}}+\lg{nm})$ tiles.
Let $\mathcal{T}_{P}$ be a TAS that strictly self assembles $P$ using the skeleton construction from 4.1. We will use this assembly as a sub component of the larger assembly that builds $GRID\_REPEAT(P,m)$. We divide the skeleton into $\frac{n}{\log n}$ parts called spines. Each spine is the vertical bar that extends upwards (the shaft) and a subsection of bottom row of the skeleton extending as far as the ribs that connect to the spine (the base). The skeleton construction will also be modified to use temp 2 and have the bottom row extend the length of the full square. This does not change the tile complexity of the construction.
Now we present the construction. Given a pattern $P$ and an integer $m$ a TAS $\mathcal{T}_{R}$ is built. Let $T_{COUNT}$ be a constant set of counter tiles. For each tile in $T_{COUNT}$ we create a copy of each spine from $\mathcal{T}_{P}$. The new spine will have the glues from the counter tile appended to the north of the shaft and the south, east, and west ends of the base. A copy of the rib tiles will also be made that has the counter tile’s north glue appended to their glue. This allows for the entire spine to be constructed and other spines to grow off of it according to the growth of the counter tiles. We repeat this process again to create a version of $\mathcal{T}_{P}$ where the structure is rotated 90 degrees so it grows to the east, but the pattern is still facing upwards. This will grow to the east end of the square.
Then another copy of the tiles from $\mathcal{T}_{P}$ is added to fill in the rest of the square with the pattern. There are a constant number of copies of $\mathcal{T}_{P}$ in the assembly so the tile complexity of these sections is the tile complexity of building any square pattern. The 4.1 theorem showed this to be $O(\frac{n^{2}}{\log n})$.
The seed of $\mathcal{T}_{R}$ will embed the starting value of the counter such that it counts $m$ times before it stops. This means that $\log m$ spines will be needed to encode the starting value. Only the bases of the spines will need to be hard coded in the seed. The bases are $\log{n}$ tiles long, so the seed will need to be a hard coded $\log nm$ tile types. If $\log m>\frac{n}{\log n}$ then the seed will span multiple copies of $P$. The seed will also travel upwards to set the starting counter value for the rotated section of the assembly.
Correctness of construction
The construction outputs a TAS $\mathcal{T}_{R}$ that is designed to strictly self-assemble $GRID\_REPEAT(P,m)$ for all patterns and values of $m$. It has a seed of a single tile and grows the entire skeleton from it. Our prior analysis shows that the tile complexity is correct at $O(\frac{n^{2}}{\log n}+\log{nm})$. Thus 5.1 is proved.
Figure 5: An example of an assembly that repeats a pattern $m=5$ times horizontally and vertically. Each spine is colored solely for clarity of presentation and in the actual construction the colors of the tiles on the spines would match the pixels of the pattern. Red spines represent a 1 and blue spines represent a 0. The spines count upwards until the counter is finished
Figure 6: An example of how the binding of a spine works. The arrows indicate the growth of the tiles. The letter indicate glues. In this example, after the blue spine at the bottom finishes growing it can grow the next red spine. The incoming blue spine tiles cooperatively binds with this red spine and the ribs from the spine below. This tile will grow the rest of the spine and the pattern continues.
6 Multilayered patters
In this section, we prove that there exist patterns, both finite and infinite, that can be weakly self-assembled using asymptotically fewer tile types by barely-3DaTAM systems than by regular, 2D aTAM systems.

### 10 Mar 2024 v2

@@ 4 Tight Bounds for Patterns on $n\times n$ Squares
In this section, we prove tight bounds on the tile complexity of self-assembling 2-colored patterns on the surfaces of $n\times n$ squares for almost all such patterns.
@@
For almost all positive integers $n$ and $P\in\texttt{SQPATS}_{2,n}$, the tile complexity of weakly self-assembling $P$ by a singly-seeded system in the aTAM is $\Theta\left(\frac{n^{2}}{\log n}\right)$.
@@
For almost all patterns $P\in\texttt{SQPATS}_{2}$, the tile complexity of weakly self-assembling $P$ by a singly-seeded system in the aTAM is $\Omega\left(\frac{n^{2}}{\log n}\right)$.
The proof of Lemma 1 is a straight-forward information-theoretic argument and can be found in Section 8.
To prove the upper bound for Theorem 4.1, we prove the following, which is a stronger result that applies to all positive integers $n$.
(b) The square once the ribs of the skeleton have filled in (blue growing to the left, yellow growing to the right).
Figure 4: A schematic example of the construction of the proof of Lemma 2. Instead of showing the black and white colors corresponding to the pattern, we color the tiles to show the pieces of the construction to which they belong.
Lemma 2
For all positive integers $n$ and $P\in\texttt{SQPATS}_{2,n}$, there exists an aTAM system $\mathcal{T}=(T,\sigma,1)$ such that $|\sigma|=1$, $|T|=O\left(\frac{n^{2}}{\log n}\right)$ and $\mathcal{T}$ weakly self-assembles $P$.
@@
We proceed by construction. Let $n\in\mathbb{Z}^{+}$ be the dimensions of the square and $P$ be the $n\times n$ pattern of black and white pixels to weakly self-assemble on the square. Our construction will yield a system $\mathcal{T}=(T,\sigma,1)$ that self-assembles an $n\times n$ square on which $P$ is formed by the black and white tiles of $T$. The tile set $T$ will be composed of two subsets, $T_{s}$ whose tiles form the skeleton, and $T_{r}$ whose tiles form the ribs. We first explain the formation of the skeleton, then that of the ribs. Figure 4 shows a high-level depiction.
@@
Skeleton The seed is part of the skeleton and is placed at location $(\lfloor\log n\rfloor,0)$ and is given the color $\texttt{Color}(P,(\lfloor\log n\rfloor,0))$. Since a vertical column of the skeleton has width one, and the ribs growing off of each side have length $\lfloor\log n\rfloor$, the width of a pair of ribs and its skeleton column (which we will call a rib-pair) is $12\lfloor\log n\rfloor+$ 1. Dividing the full width $n$ by the width of a rib-pair, and taking the floor, gives the number of full rib-pairs that will fit. Let $f=\lfloor\frac{n}{2\log n+1}\rfloor$ be this number. Let $r=n\mod(2\lfloor\log n\rfloor+1)$ be the remaining width after the last full rib-pair. If $r<\lfloor\log n\rfloor+1$, then a column of the skeleton grows up immediately to the right of the last full rib-pair, and its ribs are of length $r-1$ and grow to the right. If $r\geq\lfloor\log n\rfloor+1$, then the last skeleton column grows upward $\lfloor\log n\rfloor$ positions to the right of the last full rib-pair and has full-length ribs (i.e., $\lfloor\log n\rfloor$) that grow to its left and ribs of length $r-(\lfloor\log n\rfloor+1)$ grow to its right. In the first case, the row of the skeleton that forms the bottom row of the square extends from the seed to $x$-coordinate $f(2\lfloor\log n\rfloor+1)+1$. In the second case, that row extends from the seed to $x$-coordinate $f(2\lfloor\log n\rfloor+1)+\lfloor\log n\rfloor+1$. The tiles of that row are hard-coded and there are $O(n)$ of them. Starting with the seed and then occurring at every $12\lfloor\log n\rfloor+$ 1 locations of the bottom row, a hard-coded set of tiles grows a column of height $n-1$. This row and set of columns are the full skeleton. The number of tile types is $O(n)$ for the row and $O(n)$ for each of the $O\left(\frac{n}{\log n}\right)$ columns, for a total of $O(n)+O\left(\frac{n^{2}}{\log n}\right)=O\left(\frac{n^{2}}{\log n}\right)$ tile types. Note that each skeleton tile type is given the color of the corresponding location in the pattern $P$.
@@
Ribs From the east and west sides of each location on the columns of the skeleton, ribs grow. Each rib is composed of $\lfloor\log n\rfloor$ tiles (except the ribs growing from the easternmost column, which may be shorter). Since there are two possible colors for each of the $\lfloor\log n\rfloor$ locations of a rib, there are a maximum of $2^{\lfloor\log n\rfloor}\leq n$ possible color patterns for any rib to match the corresponding locations in $P$. (Note that we will discuss the construction of the tiles for ribs that grow to the east, and for ribs that grow to the west the directions are simply reversed.) For any given rib $r$, let the portion of $P$ corresponding to the locations of $r$ be represented by the binary string of length $\lfloor\log n\rfloor$ where each black location is represented by a 00, and each white by a $1$. For example, for a rib $r$ of length $5$ growing eastward from a column, if the corresponding locations of $P$ are “black, black, white, black, white”, then the binary string will be “$00101$”. For each possible binary string $b$ of length $\lfloor\log n\rfloor$, i.e. $b\in\{0,1\}^{\lfloor\log n\rfloor}$, a unique tile type, $t_{b}$, is made. with the glue $b$ on its west side and the glue $b[1:]$ (i.e. $b$ with its left bit truncated) on its east side. This tile type is given the color corresponding to the first bit of $b$. Additionally, for each skeleton column tile from which a rib should grow to the east with pattern $b$, the glue $b$ will be on its east side, allowing $t_{b}$ to attach. This results in the creation of a maximum of $n$ unique tile types (and there will be another $n$ for the first tiles of each westward growing rib). Recall that the tile types for the skeleton were already accounted for and each is hard-coded so that the placement of these glues does not require any new tile types for the skeleton.
@@
Now, the process is repeated for each binary string from length $b-1$ to $1$, with the color of each tile being set to the value of the first remaining bit. Each iteration requires half as many tile types to be created as the previous, i.e. $2^{\lfloor\log n\rfloor-1}$, then $2^{\lfloor\log n\rfloor-2}$, $\ldots$, $2$. Intuitively, each rib position has glues that encode their bit value in the pattern and the portion of the pattern that must be extended outward from them, away from the skeleton. Therefore, for the last position on the tip of each rib, there are exactly 2 choices, white or black, and so all ribs share from a set of two tile types made specially for the ends of ribs. For the tile types of ribs that grow to the east, the total summation is $\Sigma^{\log n-1}_{x=0}{2^{\log n-x}}=2n-2=O(n)$. Accounting for the additional tile types needed for westward growing ribs, the full tile complexity of the ribs is $O(n)$.
@@
Correctness of construction The system $\mathcal{T}$ designed to weakly self-assemble $P$, as discussed, has a seed of a single tile, and since all tile attachments require forming a bond with a single neighbor, the temperature of the system can be $\tau=1$. Our prior analysis shows that the tile complexity is correct at $O\left(\frac{n^{2}}{\log n}\right)$, and showing that $\mathcal{T}$ weakly self-assembles $P$ is trivial since (1) the tiles of the skeleton are specifically hard-coded to be colored for their corresponding locations in $P$, and (2) for each possible pattern corresponding to a rib there is a hard-coded set of rib tiles that match that pattern and grow from the skeleton into those locations. Thus, $P$ is formed and Lemma 2 is proved, and with both Lemmas 1 and 2, Theorem 4.1 is proved. (Example aTAM systems for this construction, as well as software capable of generating other systems for patterns derived from image files, and for simulating them, can be found online [7].)
@@ 5 Repeated Patterns
Figure 5: An example of an assembly that repeats a pattern $m=5$ times horizontally and vertically. Each spine is colored solely for clarity of presentation, and in the actual construction, the colors of the tiles on the spines would match the pixels of the pattern. Red spines represent a 1, and blue spines represent a 0. The spines count upwards until the counter is finished
In this section, we discuss patterns consisting of repeated, arbitrary square sub-patterns, and that efficient systems exist that weakly self-assemble them.
@@
Given $n,m\in\mathbb{Z}^{+}$ and $P\in\texttt{SQPATS}_{2,n}$, define $\texttt{GridRepeat}(P,m)\rightarrow P^{\prime}$ such that $P^{\prime}\in\texttt{SQPATS}_{2,nm}$ is an $nm\times nm$ square consisting of an $m\times m$ square composed of an $n\times n$ grid of copies of the pattern $P$.
@@
For all $n,m\in\mathbb{N}$, and $P\in\texttt{SQPATS}_{2,n}$ there exists an aTAM system $\mathcal{T}=(T,\sigma,2)$ such that $|\sigma|=1$, $|T|=O(\frac{n^{2}}{\log{n}}+\log{mn})$ and $\mathcal{T}$ weakly self-assembles $\texttt{GridRepeat}(P,m)$.
The proof of Theorem 5.1 can be found in Section 9. It makes use of an extension of the construction for the proof of Lemma 2 and embeds a counter into the skeleton and ribs so that the copies of the sub-pattern are correctly counted.
6 Barely-3DaTAM patterns
In this section, we show that there exist patterns, both finite and infinite, that can be weakly self-assembled using exponentially fewer tile types by barely-3DaTAM systems than by any regular, 2D aTAM systems. (We also note that the exponential separation can be increased arbitrarily.)